# Daniel Iong, Matthew McAnear, Yuezhou Qu

### v1

@@ Abstract
Keywords Guassian Process  $\cdot$ regression  $\cdot$ contaminated normal  $\cdot$ SuperMAG  $\cdot$ DeltaB
@@ 1 Introduction
In this context, we consider geomagnetic perturbation measurements from various ground magnetometer stations around the globe from 2010 to 2015 that we obtained from SuperMAG [1]. This data provides a proxy for measuring geomagnetically induced currents, which could potentially drive catastrophic disruptions to critical infrastructure, such as power grids and oil pipelines [2]. Therefore, it is imperative that we are able to obtain accurate predictions and predictive uncertainty estimates for these quantities.
Our goal is to predict the maximum value of the horizontal magnetic perturbation in the north-south direction, denoted by $\delta B_{H}$, over twenty minute intervals across twelve test stations. This data is characterized by occasional large spikes that occur during geomagnetic storms followed by long periods of relatively low variance. In this context, the influence of outliers is especially pronounced. We plot $\delta B_{H}$ in 2013 in fig. 1 as an illustrative example. We use solar wind and interplanetary magnetic field measurements obtained from NASA’s OMNIWeb [3] as drivers.
@@
The CN distribution is a special case of a Gaussian mixture distribution with two components [4]. This distribution models outliers explicitly by assigning them to a mixture component with much larger variance. Therefore, the mixture proportions can be interpreted as the outlier and non-outlier proportion. Robust regression with mixture noise distributions is not a new concept. [5] first introduced a Bayesian linear regression model with mixture noise as a Bayesian approach to handling outliers. [6] follows up with a variational approximation for a similar model. Although the idea of modeling noise with mixture distributions in regression dates back to the 1960s, it is still being actively explored as a means to robustify modern statistical and machine learning methods. [7] proposed DeepGUM, a deep regression model with Gaussian-uniform mixture noise for performing various computer vision tasks in the presence of outliers. [8] developed the Mixture-of-Gaussian Lasso technique which models noise as a Gaussian mixture distribution in a sparse linear regression model. [9] introduced a probabilistic principal component regression model with switching Gaussian mixture noise for industrial process modeling.
@@
GPR with mixture noise was first proposed by [10] who developed an expectation propagation (EP) algorithm and Markov chain Monte-Carlo (MCMC) sampling scheme for inference. [10] also developed inference algorithms for GPR models with Student-t and Laplace noise. Their experiments showed that the robust GPR models outperform the GPR model with Gaussian noise when outliers are present in the training data. However, predictive performance among the robust GPR models were similar for the simulated and real-world datasets they considered. [11] also considered a GPR model with mixture noise but used an EM algorithm for inference. [12] proposed the twinned GP model which also assumes a mixture noise distribution. However, they model the outlier proportion with another GP, effectively making their noise model heteroscedastic which is out of the scope of our proposed work. They showed that the twinned GP model is suitable in cases where outliers can be clustered. A major issue with the inference algorithms developed by [10], [11], and [12] is that they scale poorly with number of observations as they perform inference on exact GPs which require inverting a full kernel matrix.
@@
In addition to GPR models with mixture noise, several other robust GPR methods have also been proposed. [13] provided a robust expectation propagation (EP) algorithm for GPR with student-t noise. [14] proposed iterative trimming for robust GPR. The main idea of their method is to iteratively trim a proportion of the observations with the largest absolute residuals so that they are not as influential to the resulting model fit. However, this may not be appropriate in cases where outliers are due to systematic effects not captured by the model but are still of interest to the analysis. [15] recently proposed using a Huber likelihood for robust GPR. This method employs weights based on projection statistics to scale residuals and bound the influence of outliers on the latent function estimate. Unfortunately, all of these methods also perform exact inference.
@@
Many approximation methods for scaling up GPR to accomodate massive datasets have been proposed [16]. These approximation methods can be categorized as either global or local approximations. Global approximations approximate the kernel matrix, and as a result the latent function posterior, through global distillation. This can be achieved by computing the kernel matrix on only a subset of the training data, sparsifying the kernel matrix, or constructing a sparse approximation of the latent function posterior using a small number of inducing (or pseudo) points. Local approximations take a divide and conquer approach by focusing on local subsets of training data. While global approximations are better at capturing global patterns, they often filter out local patterns due to the use of inducing points. On the other hand, local approximations are better at capturing non-stationary features but risk local over-fitting. [16] provides a comprehensive review of both global and local approximation methods for GPR.
@@
Sparse approximations have perhaps received the most attention in the GP community. Early work in this area involved approximating the GP prior with inducing points and then optimizing the marginal likelihood of the approximate model [17]. This includes the deterministic training conditional (DTC), fully independent training conditional (FITC), and partially independent training conditional (PITC) approximations which only differ in how they specify the dependency between the latent function and inducing points [18, 19, 20]. [21] further extended these approximated GP priors to be more flexible by placing the inducing points in a different domain via integral transforms. However, these approaches effectively alter the original model assumptions and may be prone to overfitting. This motivated the development of approximate inference approaches that make all the necessary approximations at inference time.
@@
[22] first introduced the sparse variational GP (SVGP) method. A subtle but important distinction between this approach and earlier approaches is that the inducing points are now considered variational parameters and are decoupled from the original model. [22] jointly estimates the inducing points and model hyperparameters by maximizing a lower bound to the exact marginal likelihood as opposed to maximizing an approximate marginal likelihood. [23] extended this idea by reformulating the lower bound to enable stochastic optimization. [24] further modified this lower bound to accomodate non-Gaussian likelihoods and applied it to GP classification. Sparse variational methods are discussed in more detail in section 2.2. In this paper, we propose a scalable inference algorithm based on the SVGP method for fitting sparse GPR models with contaminated normal (CN) noise on large datasets.
@@
The rest of the paper is organized as follows. Section 2 provides background on Gaussian process regression and sparse variational GPs. We then discuss our proposed model in section 3.1 followed by a corresponding inference algorithm in section 3.2. In section 4.1, we perform a simulation study to show the efficacy of our proposed inference algorithm. In section 4.2, we compare sparse GPR models with different noise distributions (CN, Gaussian, Student-t, Laplace) trained on simulated datasets. In section 4, we train sparse GPR models with different noise distributions on flight delays and ground magnetic perturbations data and compare their predictive performance. We conclude the paper with a brief summary and discussion on potential extensions in section 6. For all simulation studies and real-world data applications, we use the GPyTorch Python package to train GPR models [25].
@@ 2.1 Gaussian process regression
A more comprehensive review of Gaussian process regression can be found in [26]. Computing eq. 5 involves inverting the $n\times n$ matrix $\tilde{\mathbf{K}}_{nn}$ which requires $O(n^{3})$ computations. This restricts the use of exact GPR models to datasets with up to a few thousand observations. Sparse approximations can be used to make GPR more scalable. Sparse GP methods approximate the latent function posterior by conditioning on a small number $m<<n$ of inducing points that act as a representative proxy for the observed outputs. These inducing points can either be chosen as a subset of the training data or optimized over. In the remainder of this section, we will discuss popular methods for jointly estimating the inducing points and hyperparameters.
@@
Note that integrating out $\mathbf{u}$ returns us to the original model so they are equivalent in terms of performing inference. However, inference with this augmented model still requires inverting $n\times n$ matrices. Several methods have been proposed to approximate the distribution in eq. 6, thereby modifying the GP prior and likelihood [17]. The fully independent training conditional (FITC) approximation removes the conditional dependencies between different elements in $\mathbf{f}$, i.e.
@@
where $\tilde{\mathbf{Q}}_{nn}=\mathbf{Q}_{nn}+\text{diag}(\mathbf{K}_{nn}-\mathbf{Q}_{nn})$ is an approximation to the true covariance $\mathbf{K}_{nn}$. Similarly, the deterministic training conditional (DTC) and partially independent training conditional (PITC) methods approximate $\mathbf{K}_{nn}$ with $\mathbf{Q}_{nn}$ and a block diagonalization of $\mathbf{K}_{nn}-\mathbf{Q}_{nn}$, respectively. The inducing points and hyperparameters can then be jointly estimated by maximizing the marginal log likelihood. With these approximations, the cost of inference and prediction is reduced from $O(n^{3})$ to $O(nm^{2})$. However, these methods are philosophically troubling as they entangle assumptions about the data embedded in the original likelihood with the approximations required to perform inference. Furthermore, $m$ new model parameters are added which increases the risk of overfitting. The sparse variational GP (SVGP) method, first introduced by [22], takes a different approach by approximating the exact posterior GP with variational inference. Before discussing this approach, we take a detour to give the reader a primer on variational inference.
@@ 2.2 Sparse variational GPs (SVGP)
where $\tilde{\mathbf{K}}_{nn}=\mathbf{K}_{nn}-\mathbf{Q}_{nn}$; and $\text{Tr}(\mathbf{A})$ denotes the trace of matrix $\mathbf{A}$. The derivation for this lower bound is given in appendix B. In order to maximize the ELBO using stochastic variational inference, [23] proposed maintaining an explicit variational distribution given by
@@
$\sum_{i=1}^{n}\Big{[}\log N(y_{i}|\mathbf{k}_{i}^{\prime}\mathbf{K}_{mm}^{-1}\mathbf{m},\sigma^{2})-\frac{1}{2}\sigma^{-2}\tilde{\mathbf{k}}_{ii}-\frac{1}{2}\text{Tr}(\mathbf{S}\mathbf{\Lambda}_{i})\Big{]}-\text{KL}[q(\mathbf{u})||p(\mathbf{u})],$
where $\mathbf{k}_{i}$ is the $i$th column of $\mathbf{K}_{mn}$; $\tilde{\mathbf{k}}_{ii}$ is the $i$th diagonal element of $\tilde{\mathbf{K}}_{nn}$; and $\mathbf{\Lambda}_{i}=\sigma^{-2}\mathbf{K}_{mm}^{-1}\mathbf{k}_{i}\mathbf{k}_{i}^{\prime}\mathbf{K}_{mm}^{-1}$ (Section 3.1 of [23]). Furthermore, the approximate predictive distribution at a new input $\mathbf{x}^{*}$ is given by
@@ 3.2 Inference
Following the same logic in appendix A, the log likelihood function $\log p(y_{i}|f_{i},\theta)$ can be decomposed as
@@
$=\sum_{i=1}^{n}\Big{\{}E_{q(f_{i},\alpha_{i}|\varphi,\vartheta)}\big{[}\log p(y_{i},\alpha_{i}|f_{i},\theta)\big{]}+$ (17)
$E_{q(f_{i}|\varphi,\vartheta)}\big{[}H(q(\alpha_{i}|f_{i}))\big{]}\Big{\}}-\text{KL}[q(\mathbf{u})||p(\mathbf{u});\vartheta]$
@@
where $\Theta=(\theta,\vartheta)$ denotes the model and kernel hyperparameters and $q(f_{i},\alpha_{i}|\varphi,\vartheta)=q(\alpha_{i}|f_{i})q(f_{i}|\varphi,\vartheta)$. We assume $q(f_{i}|\varphi,\vartheta)$ is the Gaussian variational distribution given in eq. 13. Although the true latent function posterior may have multiple modes in this case, [10] showed that a Laplace approximation for the latent function posterior of an exact GP with mixture noise works reasonably well in the presence of a few outliers. Since $q(\textbf{u})$ and $p(\textbf{u})$ are both Gaussian, the KL divergence term can be expressed as a sum of $n$ terms and therefore, $\mathcal{L}(q,\Theta)$ can be written as $\mathcal{L}(q,\Theta)=\sum_{i=1}^{n}\mathcal{L}_{i}(q,\Theta)$. This allows it to be maximized using stochastic optimization methods. For the rest of this section, we describe the SGAM algorithm for maximizing the modified ELBO (or equivalently, minimizing the negative modified ELBO).
@@
where $\tilde{\Theta}^{(t)}=(\theta^{(t)},\vartheta)$; and $\alpha_{2}^{(t)}$ is the learning rate at iteration $t$. To make this more computationally efficient, we can also take a SGD step with respect to $\theta$ to obtain $\theta^{(t)}$ instead of computing the closed-form updates in the first part of the backward step. Note that the $\tau^{(t)}$ update in eq. 21 is not constrained to be greater than 1. If the final $\tau$ estimate is less than 1, we can replace $\sigma^{2}$ with $\tau\sigma^{2}$, $\tau$ with $\tau^{-1}$, and $\pi$ with $1-\pi$ so that $\tau$ and $\pi$ can still be interpreted as the variance inflation parameter and outlier probability, respectively. The approximate posterior predictive distribution at a new input $\mathbf{x}^{*}$ is given by
$q_{\text{pred}}(y^{*}|\mathbf{y},\mathbf{x}^{*})=\pi N(y^{*}|\mu_{f^{*}},\sigma_{f^{*}}^{2}+\tau\sigma^{2})+(1-\pi)N(y^{*}|\mu_{f^{*}},\sigma^{2}),$ (22)

### v3

# Sparse Variational Contaminated Noise Gaussian Process Regression with Applications in Geomagnetic Perturbations Forecasting

@@ Abstract
keywords:
Guassian Process; regression; contaminated normal; SuperMAG; DeltaB;
@@ 1 Introduction
In this context, we consider geomagnetic perturbation measurements from various ground magnetometer stations around the globe from 2010 to 2015 that we obtained from SuperMAG (Gjerloev, 2012). This data provides a proxy for measuring geomagnetically induced currents, which could potentially drive catastrophic disruptions to critical infrastructure, such as power grids and oil pipelines (Schrijver et al., 2014). Therefore, it is imperative that we are able to obtain accurate predictions and predictive uncertainty estimates for these quantities.
Our goal is to predict the maximum value of the horizontal magnetic perturbation in the north-south direction, denoted by $\delta B_{H}$, over twenty minute intervals across twelve test stations. This data is characterized by occasional large spikes that occur during geomagnetic storms followed by long periods of relatively low variance. In this context, the influence of outliers is especially pronounced. We plot $\delta B_{H}$ in 2013 in fig. 1 as an illustrative example. We use solar wind and interplanetary magnetic field measurements obtained from NASA’s OMNIWeb (Papitashvili and King, 2020) as drivers of $\delta B$ (table 1).
Feature Description Units
$\text{bx}_{gse}$ X component of magnetic vector in GSE coordinates nT
$\text{by}_{gse}$ Y component of magnetic vector in GSE coordinates nT
$\text{bz}_{gse}$ Z component of magnetic vector in GSE coordinates nT
v Solar Wind Speed km/s
density Proton Density $N/cm^{3}$
temperature Proton Temperature K
pressure Flow Pressure nPa
e Electric Field mV/m
Table 1: Features used in prediction of $\delta$ B.
@@
The CN distribution is a special case of a Gaussian mixture distribution with two components (Gleason, 1993). This distribution models outliers explicitly by assigning them to a mixture component with much larger variance. Therefore, the mixture proportions can be interpreted as the outlier and non-outlier proportion. Robust regression with mixture noise distributions is not a new concept. Box and Tiao, (1968) first introduced a Bayesian linear regression model with mixture noise as a Bayesian approach to handling outliers. Faul and Tipping, (2001) follows up with a variational approximation for a similar model. Although the idea of modeling noise with mixture distributions in regression dates back to the 1960s, it is still being actively explored as a means to robustify modern statistical and machine learning methods. Lathuilière et al., (2018) proposed DeepGUM, a deep regression model with Gaussian-uniform mixture noise for performing various computer vision tasks in the presence of outliers. Xu and Zhang, (2019) developed the Mixture-of-Gaussian Lasso technique which models noise as a Gaussian mixture distribution in a sparse linear regression model. Sadeghian et al., (2022) introduced a probabilistic principal component regression model with switching Gaussian mixture noise for industrial process modeling.
@@
GPR with mixture noise was first proposed by Kuss, (2006) who developed an expectation propagation (EP) algorithm and Markov chain Monte-Carlo (MCMC) sampling scheme for inference. Kuss, (2006) also developed inference algorithms for GPR models with Student-t and Laplace noise. Their experiments showed that the robust GPR models outperform the GPR model with Gaussian noise when outliers are present in the training data. However, predictive performance among the robust GPR models were similar for the simulated and real-world datasets they considered. Daemi et al., (2019) also considered a GPR model with mixture noise but used an EM algorithm for inference. Naish and Holden, (2007) proposed the twinned GP model which also assumes a mixture noise distribution. However, they model the outlier proportion with another GP, effectively making their noise model heteroscedastic which is out of the scope of our proposed work. They showed that the twinned GP model is suitable in cases where outliers can be clustered. A major issue with the inference algorithms developed by Kuss, (2006), Daemi et al., (2019), and Naish and Holden, (2007) is that they scale poorly with number of observations as they perform inference on exact GPs which require inverting a full kernel matrix.
@@
In addition to GPR models with mixture noise, several other robust GPR methods have also been proposed. Jylänki et al., (2011) provided a robust expectation propagation (EP) algorithm for GPR with student-t noise. Li et al., (2021) proposed iterative trimming for robust GPR. The main idea of their method is to iteratively trim a proportion of the observations with the largest absolute residuals so that they are not as influential to the resulting model fit. However, this may not be appropriate in cases where outliers are due to systematic effects not captured by the model but are still of interest to the analysis. Algikar and Mili, (2023) recently proposed using a Huber likelihood for robust GPR. This method employs weights based on projection statistics to scale residuals and bound the influence of outliers on the latent function estimate. Gu et al., (2018) discusses methods for robustly estimating the covariance parameters in a GP model with Gaussian noise. Unfortunately, all of these methods also perform exact inference. Altamirano et al., (2023) proposes a robust and conjugate Gaussian process that also performs exact inference but can be plugged into scalable GP approximations. Their method involves doing generalized Bayesian inference with a robust loss function whereas our proposed method involves an explicit robust noise assumption.
@@
Many approximation methods for scaling up GPR to accomodate massive datasets have been proposed (Liu et al., 2019). These approximation methods can be categorized as either global or local approximations. Global approximations approximate the kernel matrix, and as a result the latent function posterior, through global distillation. This can be achieved by computing the kernel matrix on only a subset of the training data, sparsifying the kernel matrix, or constructing a sparse approximation of the latent function posterior using a small number of inducing (or pseudo) points. Local approximations take a divide and conquer approach by focusing on local subsets of training data. While global approximations are better at capturing global patterns, they often filter out local patterns due to the use of inducing points. On the other hand, local approximations are better at capturing non-stationary features but risk local over-fitting. Liu et al., (2019) provides a comprehensive review of both global and local approximation methods for GPR.
@@
Sparse approximations have perhaps received the most attention in the GP community. Early work in this area involved approximating the GP prior with inducing points and then optimizing the marginal likelihood of the approximate model (Quiñonero-Candela and Rasmussen, 2005). This includes the deterministic training conditional (DTC), fully independent training conditional (FITC), and partially independent training conditional (PITC) approximations which only differ in how they specify the dependency between the latent function and inducing points (Seeger et al., 2003; Snelson and Ghahramani, 2005, 2007). Lázaro-Gredilla and Figueiras-Vidal, (2009) further extended these approximated GP priors to be more flexible by placing the inducing points in a different domain via integral transforms. However, these approaches effectively alter the original model assumptions and may be prone to overfitting. This motivated the development of approximate inference approaches that make all the necessary approximations at inference time.
@@
Titsias, (2009) first introduced the sparse variational GP (SVGP) method. A subtle but important distinction between this approach and earlier approaches is that the inducing points are now considered variational parameters and are decoupled from the original model. Titsias, (2009) jointly estimates the inducing points and model hyperparameters by maximizing a lower bound to the exact marginal likelihood as opposed to maximizing an approximate marginal likelihood. Hensman et al., (2013) extended this idea by reformulating the lower bound to enable stochastic optimization. Hensman et al., (2015) further modified this lower bound to accomodate non-Gaussian likelihoods and applied it to GP classification. Sparse variational methods are discussed in more detail in section 2.2. In this paper, we propose a scalable inference algorithm based on the SVGP method for fitting sparse GPR models with contaminated normal (CN) noise on large datasets.
@@
The rest of the paper is organized as follows. Section 2 provides background on Gaussian process regression and sparse variational GPs. We then discuss our proposed model in section 3.1 followed by a corresponding inference algorithm in section 3.2. In section 4.1, we perform a simulation study to show the efficacy of our proposed inference algorithm. In section 4.2, we compare sparse GPR models with different noise distributions (CN, Gaussian, Student-t, Laplace) trained on simulated datasets. In section Section 5, we train GPR models with variable noise distributions on flight delays and ground magnetic perturbations data and compare their predictive performance. We conclude the paper with a brief summary and discussion on potential extensions in section 6. For all simulation studies and real-world data applications, we use the GPyTorch Python package to train GPR models (Gardner et al., 2018).
@@ 2.1 Gaussian process regression
A more comprehensive review of Gaussian process regression can be found in Rasmussen and Williams, (2005). Computing eq. 5 involves inverting the $n\times n$ matrix $\tilde{\mathbf{K}}_{nn}$ which requires $O(n^{3})$ computations. This restricts the use of exact GPR models to datasets with up to a few thousand observations. Sparse approximations can be used to make GPR more scalable. Sparse GP methods approximate the latent function posterior by conditioning on a small number $m<<n$ of inducing points that act as a representative proxy for the observed outputs. These inducing points can either be chosen as a subset of the training data or optimized over. In the remainder of this section, we will discuss popular methods for jointly estimating the inducing points and hyperparameters.
@@
Note that integrating out $\mathbf{u}$ returns us to the original model so they are equivalent in terms of performing inference. However, inference with this augmented model still requires inverting $n\times n$ matrices. Several methods have been proposed to approximate the distribution in eq. 6, thereby modifying the GP prior and likelihood (Quiñonero-Candela and Rasmussen, 2005). The fully independent training conditional (FITC) approximation removes the conditional dependencies between different elements in $\mathbf{f}$, i.e.
@@
where $\tilde{\mathbf{Q}}_{nn}=\mathbf{Q}_{nn}+\text{diag}(\mathbf{K}_{nn}-\mathbf{Q}_{nn})$ is an approximation to the true covariance $\mathbf{K}_{nn}$. Similarly, the deterministic training conditional (DTC) and partially independent training conditional (PITC) methods approximate $\mathbf{K}_{nn}$ with $\mathbf{Q}_{nn}$ and a block diagonalization of $\mathbf{K}_{nn}-\mathbf{Q}_{nn}$, respectively. The inducing points and hyperparameters can then be jointly estimated by maximizing the marginal log-likelihood. With these approximations, the cost of inference and prediction is reduced from $O(n^{3})$ to $O(nm^{2})$. However, these methods are philosophically troubling as they entangle assumptions about the data embedded in the original likelihood with the approximations required to perform inference. Furthermore, $m$ new model parameters are added which increases the risk of overfitting. The sparse variational GP (SVGP) method, first introduced by Titsias, (2009), takes a different approach by approximating the exact posterior GP with variational inference. Before discussing this approach, we take a detour to give the reader a primer on variational inference.
@@ 2.2 Sparse variational GPs (SVGP)
where $\tilde{\mathbf{K}}_{nn}=\mathbf{K}_{nn}-\mathbf{Q}_{nn}$; and $\text{Tr}(\mathbf{A})$ denotes the trace of matrix $\mathbf{A}$. The derivation for this lower bound is given in appendix B. In order to maximize the ELBO using stochastic variational inference, Hensman et al., (2013) proposed maintaining an explicit variational distribution given by
@@
$\sum_{i=1}^{n}\Big{[}\log N(y_{i}|\mathbf{k}_{i}^{\prime}\mathbf{K}_{mm}^{-1}\mathbf{m},\sigma^{2})-\frac{1}{2}\sigma^{-2}\tilde{\mathbf{k}}_{ii}-\frac{1}{2}\text{Tr}(\mathbf{S}\boldsymbol{\Lambda}_{i})\Big{]}-\text{KL}[q(\mathbf{u})||p(\mathbf{u})],$
where $\mathbf{k}_{i}$ is the $i$th column of $\mathbf{K}_{mn}$; $\tilde{\mathbf{k}}_{ii}$ is the $i$th diagonal element of $\tilde{\mathbf{K}}_{nn}$; and $\boldsymbol{\Lambda}_{i}=\sigma^{-2}\mathbf{K}_{mm}^{-1}\mathbf{k}_{i}\mathbf{k}_{i}^{\prime}\mathbf{K}_{mm}^{-1}$ (Section 3.1 of Hensman et al., (2013)). Furthermore, the approximate predictive distribution at a new input $\mathbf{x}^{*}$ is given by
@@ 3.2 Inference
Following the same logic in appendix A, the log-likelihood function $\log p(y_{i}|f_{i},\theta)$ can be decomposed as
@@
$=\sum_{i=1}^{n}\Big{\{}E_{q(f_{i},\alpha_{i}|\varphi,\vartheta)}\big{[}\log p(y_{i},\alpha_{i}|f_{i},\theta)\big{]}$ (17)
$+E_{q(f_{i}|\varphi,\vartheta)}\big{[}H(q(\alpha_{i}|f_{i}))\big{]}\Big{\}}-\text{KL}[q(\mathbf{u})||p(\mathbf{u});\vartheta]$
@@
where $\Theta=(\theta,\vartheta)$ denotes the model and kernel hyperparameters and $q(f_{i},\alpha_{i}|\varphi,\vartheta)=q(\alpha_{i}|f_{i})q(f_{i}|\varphi,\vartheta)$. We assume $q(f_{i}|\varphi,\vartheta)$ is the Gaussian variational distribution given in eq. 13. Although the true latent function posterior may have multiple modes in this case, Kuss, (2006) showed that a Laplace approximation for the latent function posterior of an exact GP with mixture noise works reasonably well in the presence of a few outliers. Since $q(\textbf{u})$ and $p(\textbf{u})$ are both Gaussian, the KL divergence term can be expressed as a sum of $n$ terms and therefore, $\mathcal{L}(q,\Theta)$ can be written as $\mathcal{L}(q,\Theta)=\sum_{i=1}^{n}\mathcal{L}_{i}(q,\Theta)$. This allows it to be maximized using stochastic optimization methods. For the rest of this section, we describe the SGAM algorithm for maximizing the modified ELBO (or equivalently, minimizing the negative modified ELBO). We summarize the algorithm in algorithm 1.
@@
where $\tilde{\Theta}^{(t)}=(\theta^{(t)},\vartheta)$; and $\alpha_{2}^{(t)}$ is the learning rate at iteration $t$, which is set to decay exponentially starting with an initial value of 0.1. To make this more computationally efficient, we can also take a SGD step with respect to $\theta$ to obtain $\theta^{(t)}$ instead of computing the closed-form updates in the first part of the backward step. Note that the $\tau^{(t)}$ update in eq. 21 is not constrained to be greater than 1. If the final $\tau$ estimate is less than 1, we can replace $\sigma^{2}$ with $\tau\sigma^{2}$, $\tau$ with $\tau^{-1}$, and $\pi$ with $1-\pi$ so that $\tau$ and $\pi$ can still be interpreted as the variance inflation parameter and outlier probability, respectively. The approximate posterior predictive distribution at a new input $\mathbf{x}^{*}$ is given by
$q_{\text{pred}}(y^{*}|\mathbf{y},\mathbf{x}^{*})=\pi N(y^{*}|\mu_{f^{*}},\sigma_{f^{*}}^{2}+\tau\sigma^{2})+(1-\pi)N(y^{*}|\mu_{f^{*}},\sigma_{f^{*}}^{2}+\sigma^{2}),$ (22)
@@
Algorithm 1 Stochastic Generalized Alternating Maximization (SGAM)
1:  Initialize $\Theta^{(0)}=(\theta^{(0)},\vartheta^{(0)})$
2:  Initialize variational distribution $q^{(0)}(f_{i},\alpha_{i})=q(\alpha_{i}|f_{i})q(f_{i}|\phi^{(0)},\vartheta^{(0)})$
3:  for $t=1,2,\ldots$ do
4:     Sample a random subset $I_{t}$ uniformly from $\{1,\ldots,n\}$
5:     Forward Step:
6:     for each $i\in I_{t}$ do
7:        Update $q^{(t)}(\alpha_{i}|f_{i})=p(\alpha_{i}|f_{i},y_{i},\Theta^{(t-1)})$
8:     end for
9:     Update $\phi^{(t)}$ using stochastic gradient descent:
$\phi^{(t)}\leftarrow\phi^{(t-1)}+\alpha_{1}^{(t)}\nabla_{\phi}L_{I_{t}}(q_{\phi}^{(t)},\Theta^{(t-1)})$
10:     Compute $q^{(t)}(f_{i},\alpha_{i})=q(f_{i}|\phi^{(t)},\vartheta^{(t-1)})p(\alpha_{i}|f_{i},y_{i},\Theta^{(t-1)})$
11:     Backward Step:
12:     Update $\Theta$ by maximizing $L(q^{(t)},\Theta)$ with respect to $\theta$ and $\vartheta$:
$\Theta^{(t)}=\arg\max_{\Theta}L(q^{(t)},\Theta)$
13:     Alternatively, use stochastic gradient descent to update $\Theta$:
$\vartheta^{(t)}\leftarrow\vartheta^{(t-1)}+\alpha_{2}^{(t)}\nabla_{\vartheta}L_{I_{t}}(q^{(t)},\Theta)$
14:  end for